OpenStudy (nilankshi):

An object of mass 15 kg is moving with a uniform velocity of 4 m s-1. what is the kinetic energy possessed by the object?

OpenStudy (nilankshi):

help mee plzz..?

OpenStudy (anonymous):

It is given by the equation, \[E = \frac{1}{2} m v ^{2}\] Where E is kinetic energy, m is mass of the object, v is the velocity of the object. Substituting the values, E= 0.5 * 15 * 4^2 =120 J

OpenStudy (nilankshi):

thzz..